### v2

# RACOON: An LLM-based Framework for Retrieval-Augmented Column Type Annotation with a Knowledge Graph

Abstract
As an important component of data exploration and integration, Column Type Annotation (CTA) aims to label columns of a table with one or more semantic types. With the recent development of Large Language Models (LLMs), researchers have started to explore the possibility of using LLMs for CTA, leveraging their strong zero-shot capabilities. In this paper, we improve on LLM-based methods for CTA by showing how to use a Knowledge Graph (KG) to augment the context information provided to the LLM. Our approach, called Racoon, combines both pre-trained parametric and non-parametric knowledge during generation to improve LLMs’ performance on CTA. Our experiments show that Racoon achieves up to a 0.21 micro F-1 improvement compared against vanilla LLM inference.
1 Introduction
In recent years, researchers and data scientists have found abundant relational tables on the Web (Cafarella et al. (2008a, b); Bhagavatula et al. (2015)). Despite their high quality, these Web Tables often miss important meta information such as column headers and relationships between columns (Suhara et al. (2022)). This information is essential for downstream tasks such as data quality control (Schelter et al. (2018)) and data discovery (Chapman et al. (2020)). A key step in recovering this meta information is Column Type Annotation (CTA), which aims to assign one or more semantic type(s) to columns in a given table. For example, the column [’UEFA Champions League’, ’Scottish Cup’] has the annotation “time.event”.
Thus, there has been increasing interest and effort recently in developing methods that can automatically assign semantic labels to columns (Deng et al. (2022); Suhara et al. (2022); Miao and Wang (2023)). Recent work (Kayali et al. (2024); Feuer et al. (2023)) argues that relying on Pre-trained Language Models (PLMs) is not sufficient because they need task-specific and dataset-specific fine-tuning with carefully labeled training dataset to achieve reasonable results. On the other hand, the recent development of Large Language Models (LLMs) (Radford et al. (2018); Brown (2020); Ouyang et al. (2022)) with pre-trained parametric memory has opened the possibility of bridging this gap by requiring minimal or even no training examples to achieve promising performance.
There has been much research effort in developing effective solutions that utilize LLMs for CTA and table related tasks (Narayan et al. (2022), Tian et al. (2024)). For CTA specifically, one line of work (Korini and Bizer (2023)) directly applies LLMs to solve CTA with promising performance. CHORUS (Kayali et al. (2024)) and Archetype (Feuer et al. (2023)) further develop other components using history and remapping to improve performances. Table-GPT (Li et al. (2024)) fine-tunes LLMs to adapt to relational data format to improve performance on various table understanding tasks.
At the core of existing methods, they directly apply LLMs without any external knowledge. Despite promising performance, LLM-based methods still encounter challenges dealing with outdated knowledge (He et al. (2022)), producing factual inaccuracies (Ji et al. (2023)), and handling domain-specific or specialized queries (Kandpal et al. (2023)). To alleviate such problems, Retrieval-Augmented Generation (RAG), which enhances LLMs’ generation by retrieving knowledge from external sources as non-parametric memory (Asai et al. (2023)), has emerged as a promising approach. Among the many external knowledge sources, Knowledge Graphs (KGs), can provide succinct yet informative knowledge to help analyze the intricate semantics of entities (Ji et al. (2021); Wang et al. (2017)). Specifically, KGs represent information in the form of entities (nodes) and relations (edges) (Wang et al. (2017)). Each edge is a factual triplet in the form of (subject, relation, object), making KG a well-structured data source. Many recent approaches have studied integrating KGs with LLMs/PLMs through RAG to improve performances on various tasks (Xu et al. (2020); Pan et al. (2024); Dehghan et al. (2024); Sun et al. (2024); Wen et al. (2024)) and saw promising performance gains compared with vanilla LLM inference. But none of the above approaches focuses on CTA. The most relevant work KGLink ( Wang et al. (2024)) incorporates an external KG with PLM during training for CTA, thus encountering the exact PLM problems we mentioned before, whereas our work aims to leverage LLMs and KG to avoid expensive training or fine-tuning over large models.
Figure 1: Vanilla way of using LLM for CTA (left) vs Racoon (right).
Inspired by these advances, we argue that we can similarly bring external non-parametric knowledge to LLMs for CTA to improve performance and propose our framework Racoon. As shown in Figure 1, compared with vanilla LLM inference on the left, Racoon on the right unifies LLM inference with a KG by performing three additional steps: namely after ingesting a table as the input, Racoon first (1) looks at the column cells’ entity mention and retrieves related information from a KG. However, this retrieved content cannot be directly applied due to its large size and noise. Thus Racoon needs to (2) process the retrieved content, and finally (3) serialize the compressed context to augment the prompt for the LLM. Note that Racoon is orthogonal to existing LLM-based methods for CTA, which enhance LLMs’ performance through additional demonstrations and response post-processing.
In summary, this paper makes the following contributions:
(1) We introduce the problem of augmenting prompts for LLMs on CTA with external knowledge from a KG and propose an end-to-end framework Racoon for it.
(2) We explore different granularity of information Racoon can retrieve from a KG with RAG. Furthermore, we provide effective post-retrieval compression and serialization methods such that given a column of cells, we can derive additional context information to augment the prompt.
(3) We conduct experiments showing that Racoon consistently outperforms vanilla LLM inference across various scenarios and retrieval methods, achieving up to a 0.21 improvement in micro F1.
2 Racoon’s Approach
We now describe the Racoon framework. The overall workflow is shown in Figure 2. Besides a CTA query, Racoon also requires an external KG and an LLM. Note that Racoon treats the KG and LLM as black-boxes and does not require any modifications, thus enjoying a plug-and-play property. For clarity of presentation, we describe how Racoon works with a CTA query on a single column, but Racoon can process CTA queries on an arbitrary number of columns.
Figure 2: Overall Racoon workflow. Racoon performs CTA in a column-by-column manner while keeping the whole table for reference. The Retriever finds matching entities in the KG based on entity mentions in column cells and retrieves all relevant knowledge from KG. Processor reduces the retrieved cell-level knowledge into an aggregated column-level context. The compressed information is passed to Augmentor, which serializes all information into natural language and composes the KG-Context to form the final KG-Augmented Prompt.
Retriever. Inside the Retriever, the Parser first parses the input table based on the query. Given a table $T$ with a column $Col_{M}$ to perform single-column CTA, the Parser selects $C$ cells from $Col_{M}$ for retrieval purposes. As of now, Racoon uses all cells, and we leave the problem of selecting the most representative cells for future work. The selected $C$ cells are treated as text strings and passed to the KG-Linker, which takes entity mentions (the text strings) as input, connects with an external KG to link each entity mention to its referent entity in the KG, and returns the IDs of the referent entities. With those entity IDs, the final step is to further retrieve all entity labels and their one-hop neighbors in the KG as the output of the Retriever. The task KG-Linker performs is defined as Entity Linking (EL), a popular task in table interpretation ( Han et al. (2011); Deng et al. (2022)). Since many EL methods are available online, Racoon allows users to implement their own KG-Linker based on the KG they decided to use. We have experimented with two KG-Linkers: a vanilla API and a SOTA EL model ( Ayoola et al. (2022)). Note that the Retriever assumes the KG-Linker can find entities in the connected KG. If it fails to link any selected cells, Racoon will fall back to using a vanilla prompt.
Processor. The role of the Processor is to compress and refine the retrieved information from the KG, ensuring it is both relevant and concise. The retrieved information takes the form of sets of nodes and edges for each selected cell in $Col_{M}$, which is too large and noisy for LLMs to digest. Thus, the Processor takes them as input and outputs their concise representations. As of now, Racoon supports processing the retrieved information into two representations, Entity-Labels and Entity-Triplets. Entity-Labels include labels of entities in the KG. The intuition is that these labels are canonical representations of entity mentions, which are often ambiguous or incomplete. For example, the entity mention ‘15 Sge’ is linked to the entity ‘15_Sagittae’ in Wikidata KG, making it easier for LLMs to understand. Although Entity-Labels can disambiguate entity mentions, they do not provide any further information about the entities such as their types. Entity-Triplets, on the other hand, include triplets in the format (subject, relation, object). In our experiments, the linked entity serves as the subject, the relation is the instance_of relation, and the object is the entity type of which the subject is an instance. For example, (15_Sagittae, instance_of, star). Note that Entity-Labels is strictly a subset of Entity-Triplets. With this representation, Racoon fully leverages the ontological relationships among entities, enabling a deeper understanding of their types. Racoon then uses statistics to count occurrences of each entity node and relationship triplet for all selected cells in $Col_{M}$ to give summary column-level information, resembling the compress stage in traditional RAG systems.
Augmentor. Finally, the Augmentor takes the original vanilla prompt and the processed representations from the Processor to create the KG-augmented prompt. It serializes the summarized column-level KG information into natural sentences and creates the KG-Context to insert into the original vanilla prompt. We show the detailed prompt in Figure 3.
Figure 3: Vanilla (left) and Racoon (middle and right) prompts. Instruction describes the CTA task in natural language, Context describes the serialized table, and Labels shows the label set. The middle prompt uses Entity-Labels and the right prompt uses Entity-Triplets.
3 Experiments
Setup. We evaluate Racoon using the GPT-3.5 model ( Ouyang et al. (2022)) and Wikidata KG ( Vrandečić and Krötzsch (2014)) on the full test set of the WikiTables-TURL-CTA benchmark ( Deng et al. (2022)) consisting of 13,025 columns extracted and annotated in a multi-label manner using 255 Freebase’s types by the TURL team from the WikiTable corpus. The dataset provides a ground truth entity ID (Wikipedia page ID) linked to each cell in the table. On average, tables in this dataset have a mean of 21 rows and 4 entity columns (columns with at least one linked cell).
Baselines. We compare Racoon against a vanilla LLM method on CTA. Notably, both CHORUS and ArcheType ( Kayali et al. (2024); Feuer et al. (2023)) applies such vanilla LLM inference.
Table 1: The results of the baseline and Racoon in micro-F1 scores
Multi-label Single-label
KG-Linker EL Baseline ETL ETT Baseline ETL ETT
Ground Truth 1.0000 0.2609 0.4023 0.4500 0.4631 0.564 0.6814
MediaWiki API 0.4018 0.2609 0.3582 0.3494 0.4631 0.4927 0.5222
ReFinED 0.6255 0.2609 0.3698 0.3678 0.4631 0.4900 0.5705
Evaluation. We first use the ground truth entity ID linked to each cell labeled in the WikiTables-TURL-CTA dataset as the output of the KG-Linker to retrieve information from the KG. This enables us to evaluate our approach assuming a perfect KG-Linker. The results in Table 1 show that both augmented prompts using Entity-Labels (ETL) and Entity-Triplets (ETT) outperform vanilla LLM inference, with Entity-Triplets achieving the best performance overall. This highlights the potential of retrieving more comprehensive information about column cells from the KG.
Second, we evaluate Racoon’s performance with different KG-Linkers: MediaWiki API (with action=websearchentities) and a SOTA entity linking model, ReFinED ( Ayoola et al. (2022)). With either KG-linker, Racoon continues to outperform the baseline, although its performance (both Entity-Labels and Entity-Triplets) drops compared with the ground-truth linker and the drop is higher for Entity-Triplets, showing that collecting extended information is not as helpful when the initial linked entity is incorrect. The drop is less with the ReFinED KG-linker, showing that SOTA KG-linkers suffice to make our approach beneficial.
Because LLMs often underperform in the multi-label dataset setting due to the large number of possible answer combinations, we further test Racoon in the single-label setting: If the model prediction falls within the ground truth label set, we set the ground truth label to be the model prediction. Otherwise, we count it as an incorrect prediction. In this setting, the performance of both the baseline and Racoon improves. Racoon with ReFinED significantly outperform the baseline, with Entity-Triplets giving the best results.
4 Conclusion
We presented a novel end-to-end LLM-based framework Racoon for Column Type Annotation (CTA) unifying vanilla LLM and Knowledge Graph (KG) augmented inference. Racoon has three components: a Retriever for retrieving knowledge from a KG; a Processor for post-processing the retrieved content; and an Augmentor for composing the final prompt. Experiments show that Racoon consistently outperforms vanilla LLM for CTA.
Acknowledgments
This project was funded in part by a gift from NEC.
Appendix A Appendix
Here we show the example prompts for the single-label setting and the multi-label setting.
A.1 Single-label
System message: Be a helpful, accurate assistant for data discovery and exploration desiged to output valid JSON in the format {‘type’: []}
User message: Consider this table given in Comma-separated Values format: “‘ Table “‘ There are a list of 255 valid types for each column: types. Your task is to choose only one type from the list to annotate the first column. Solve this task by following these steps: 1. Look at the cells in the first column of the above table. 2. Consider this information carefully: Cells in this column are instances of the following wikidata entities: human (6 cells). 3. Choose only one valid type from the given list of types. Check that the type MUST be in the list. Give the answer in valid JSON format.
A.2 Multi-label:
System message: Be a helpful, accurate assistant for data discovery and exploration desiged to output valid JSON in the format {‘type’: []}
User message: For multi-label: Consider this table given in Comma-separated Values format: “‘ Table “‘ There are a list of 255 valid types for each column: types. Your task is to choose one or multiple types from the list to annotate the first column. Solve this task by following these steps: 1. Look at the cells in the first column of the above table. 2. Consider this information carefully: Cells in this column are instances of the following wikidata entities: human (6 cells). 3. Mark each type in the given list with 0 or 1. Mark a type with 1 if it can better represent all cells of the first column. Mark a type with 0 otherwise. 4. Give a list of types that you have marked 1 in the previous step. Check that the types MUST be in the list. Give the answer in valid JSON format.